# Alexandra Ciobotaru

### 11 Mar 2026 v3

@@ 5.2 LIWC
To determine which LIWC features have the greatest influence on predictions, we used SHapley Additive exPlanations (SHAP) [26], as explained in Section 4, with the results presented in Figure 4. The importance of the top 20 features aligns with our correlation analysis and highlights the differences between low and high scores for both depression and anxiety.
@@
The LIWC analyses presented in this section demonstrate that lexical indicators can effectively capture underlying psychological patterns related to depression and anxiety. However, the most important feature used in mental health research, the pronoun ”I”, appears to be less useful for the Romanian language, likely because it is a pro-drop language. This finding is consistent with the work of [36], which also indicated that this marker does not generalize well in a sample of patients with major depression.
5.3 Emotion Analysis
@@
Using the emotion detection model described in Section 4, we predicted the emotions in all the texts of PsihoRo, and then we aggregated the results by using radar charts to visualize the overall emotional patterns. In Figure 5 we show the emotion predictions across all PsihoRo texts, and also for the pessimistic questions (1, 3, 5) and the optimistic ones (2, 4, 6).
@@
Figure 5: Emotion predictions across: all PsihoRo texts (left), pessimistic questions (middle), and optimistic questions (right).
@@
Surprise is an emotion that is completely absent from the PsihoRo data. The following emotions are more common, in order of appearance: anger, joy, fear, sadness, trust, and the most common, neutral. The emotional pattern for the more pessimistic questions is displayed in the center of Figure 5, while the emotional pattern for the more optimistic questions is displayed on the right. Emotions such as sadness and anger are on the rise on the Pessimistic Questions, while joy and trust are more prevalent on the Optimistic Questions.
@@ 5.4.1 No Risk Group
5.4.2 Depression Risk Group
@@
5.4.3 Anxiety Risk Group
@@
6 Conclusion and Future Work
@@
We analyzed the collected data using the psychologically meaningful categories from LIWC, emotion analysis and topic modeling. We observed differences in LIWC categories related to body-related words, which may indicate somatic awareness, and tentative words, which indicate uncertainty, that correlated positively with both depression and anxiety scores. A higher preoccupation with mortality was also observed in relation to depression, while anxiety scores were positively correlated with a preoccupation with family. While LIWC analyses showed that lexical indicators can effectively capture underlying psychological patterns related to depression and anxiety, we did not find any association between the linguistic marker “I” and depression or anxiety in the Romanian language, consistent with previous research [36]. The emotion analysis revealed a prevalence of sadness in the population group at risk for depression, as well as fear in the population group at risk for anxiety, which aligns with previous research [25, 3, 38]. Topic modeling showed that although some perceptions and emotions are captured in textual responses, a multi-perspective approach that goes beyond self-expressed text is necessary to fully comprehend these patterns. By offering high-quality data with precise labels for anxiety and depression risk, PsihoRo seeks to supplement this intricate research.
7 Ethical Considerations and Limitations

### 12 Mar 2026 v4

@@ 5.2.  LIWC
Surprisingly, the use of negations is negatively correlated with both PHQ-9 and GAD-7 scores, suggesting that participants with higher depression and anxiety scores tend to use fewer negations in their responses. Individuals who score higher on the PHQ-9 and GAD-7 tend to use more words in the Tentative category, which indicates uncertainty, as evidenced by correlation coefficients of .204** for PHQ-9 and .213** for GAD-7. Comparatives also show a positive correlation with the depression score (.167*), indicating comparative thinking, which is often linked to depressive rumination Nolen-Hoeksema et al. (2008). However, another marker for rumination, the high frequency of Past focus verbs, is not observed in our data. There is a positive correlation between the use of Body-related words and PHQ-9 (.200**) and GAD-7 (.180**) scores, indicating that participants with higher scores frequently use more language related to bodily sensations. This may reflect increased somatic awareness, a common trait in depression and anxiety Penninx et al. (2013). Participants with higher PHQ-9 and GAD-7 scores discuss Achievement less frequently (-.163* and -.192**, respectively), which could indicate lower motivation. In addition, the negative correlation with Leisure-related words (-.145* and -.169*, respectively) may stem from anhedonia, a common symptom of depression Grillo (2016). Words in the Death category show a significant positive correlation with depression (.138*), but not with anxiety, indicating a greater preoccupation with mortality in participants with higher depression scores. In contrast, the Family category is positively correlated with anxiety scores (r = .145*), suggesting that individuals with higher anxiety are more focused on family-related issues.
To better understand the differences between higher and lower symptom scores for depression and anxiety, we conducted an additional analysis of the LIWC categories. According to the cut-off values for both the PHQ-9 and GAD-7, we define a low PHQ-9 score as less than 10, while a score of 10 or higher is considered high. In Figure 3, the goal was to identify linguistic markers that differentiate individuals with higher vs. lower symptom scores. Statistically significant differences are indicated with an asterisk (*), as determined by the Mann-Whitney U test. Positive values indicate greater relative use among participants with higher symptom scores, while negative values indicate greater use among low-symptom participants. The analysis produced results consistent with the correlation analysis, particularly regarding Achievement and tentativeness in relation to depression and anxiety. However, for depression, we observed a decreased use of Positive emotion words and an increased usage of Sadness words, which aligns with existing literature in psychology Trifu et al. (2024). In addition, the language of participants with high GAD-7 scores showed a higher frequency of words in the Interrogatives category, indicating a tendency to question or seek clarity.
@@
To determine which LIWC features have the greatest influence on predictions, we used SHapley Additive exPlanations (SHAP) Lundberg and Lee (2017), as explained in Section 4, with the results presented in Figure 4. The importance of the top 20 features aligns with our correlation analysis and highlights the differences between low and high scores for both depression and anxiety. For predicting depression, low values in the LIWC features for Assent, Verb, and Common Adjectives, along with high values in Tentative words, contribute to the model’s positive predictions. In the case of anxiety, low values in LIWC features related to Achievement and Insight, combined with high values in Function words, Interrogatives, and Conjunctions, also have a higher contribution to a positive prediction.
@@
The LIWC analyses presented in this section demonstrate that lexical indicators can effectively capture underlying psychological patterns related to depression and anxiety. However, the most important feature used in mental health research, the pronoun “I”, appears to be less useful for the Romanian language, likely because it is a pro-drop language. This finding is consistent with the work of Trifu et al. (2024), which also indicated that this marker does not generalize well in a sample of Romanian patients with major depression.
5.3.  Emotion Analysis
@@
Using the emotion detection model described in Section 4, we predicted emotions for all PsihoRo texts, then aggregated the results using radar charts to visualize overall emotional patterns. In Figure 5, we show the emotion predictions across all PsihoRo texts, as well as for the negative-topic (pessimistic) questions (1, 3, 5) and the positive-topic (optimistic) questions (2, 4, 6).
@@
Figure 5: Emotion predictions across: all PsihoRo texts (left), negative-topic (pessimistic) questions (middle), and positive-topic (optimistic) questions (right).
@@
Surprise is an emotion that is completely absent from the PsihoRo data. The following emotions are more common, in order of appearance: anger, joy, fear, sadness, trust, and, most commonly, neutral. The emotional pattern for the more pessimistic questions is displayed in the center of Figure 5, while the emotional pattern for the more optimistic questions is displayed on the right. Emotions such as sadness and anger are on the rise on the Pessimistic Questions, while joy and trust are more prevalent on the Optimistic Questions.
@@ 5.4.1.  No Risk Group
5.4.2.  Depression Risk Group
@@
5.4.3.  Anxiety Risk Group
@@
6.  Conclusion and Future Work
@@
We analyzed the collected data using the psychologically meaningful categories from LIWC, emotion analysis and topic modeling. We observed differences in LIWC categories related to body-related words, which may indicate somatic awareness, and tentative words, which indicate uncertainty, that correlated positively with both depression and anxiety scores. A higher preoccupation with mortality was also observed in relation to depression, while anxiety scores were positively correlated with a preoccupation with family. While LIWC analyses showed that lexical indicators can effectively capture underlying psychological patterns related to depression and anxiety, we did not find any association between the linguistic marker “I” and depression or anxiety in the Romanian language, consistent with previous research Trifu et al. (2024). The emotion analysis revealed a prevalence of sadness in the population group at risk for depression, as well as fear in the population group at risk for anxiety, which aligns with previous research Low et al. (2020); Atapattu et al. (2022); Varastehnezhad et al. (2025). Topic modeling showed that although some perceptions and emotions are captured in textual responses, a multi-perspective approach that goes beyond self-expressed text is necessary to fully comprehend these patterns. By offering high-quality data with precise labels for anxiety and depression risk, PsihoRo seeks to supplement this intricate research.
Ethical Considerations and Limitations
@@ Acknowledgments
Bibliographical References
  • M. Al-Mosaiwi and T. Johnstone (2018) In an absolute state: elevated use of absolutist words is a marker specific to anxiety, depression, and suicidal ideation. Clinical psychological science 6 (4), pp. 529–542. Cited by: §4.
  • B. Alshouha, J. Serrano‐Guerrero, F. Chiclana, F. P. Romero, and J. A. Olivas (2024) BioEmoDetector: a flexible platform for detecting emotions from health narratives. pp. 101670. External Links: Document, Link Cited by: §2.
  • T. Atapattu, M. Herath, C. Elvitigala, P. de Zoysa, K. Gunawardana, M. Thilakaratne, K. de Zoysa, and K. Falkner (2022) EmoMent: an emotion annotated mental health corpus from two South Asian countries. In Proceedings of the 29th International Conference on Computational Linguistics, N. Calzolari, C. Huang, H. Kim, J. Pustejovsky, L. Wanner, K. Choi, P. Ryu, H. Chen, L. Donatelli, H. Ji, S. Kurohashi, P. Paggio, N. Xue, S. Kim, Y. Hahm, Z. He, T. K. Lee, E. Santus, F. Bond, and S. Na (Eds.), Gyeongju, Republic of Korea, pp. 6991–7001. External Links: Link Cited by: §6.
  • A. Beech, H. Fan, J. Shu, J. Oyarzun, P. Nadel, O. T. Karaman, S. Vranos, E. A. Phelps, and M. A. Kredlow (2025) Using natural language processing to identify patterns associated with depression, anxiety, and stress symptoms during the covid-19 pandemic. 376, pp. 113–121. External Links: ISSN 0165-0327, Document, Link Cited by: §2.
  • A. Benton, G. Coppersmith, and M. Dredze (2017) Ethical research protocols for social media health research. In Proceedings of the first ACL workshop on ethics in natural language processing, pp. 94–102. Cited by: §3.
  • R. L. Boyd, A. Ashokkumar, S. Seraj, and J. W. Pennebaker (2022) The development and psychometric properties of liwc-22. Austin, TX: University of Texas at Austin 10, pp. 1–47. Cited by: §2, §4.
  • A. Briciu and M. Lupea (2018) Studying the language of mental illness in romanian social media. In 2018 IEEE 14th International Conference on Intelligent Computer Communication and Processing (ICCP), pp. 21–28. Cited by: §1.
  • A. Bucur, A. Moldovan, K. Parvatikar, M. Zampieri, A. R. KhudaBukhsh, and L. P. Dinu (2025) On the state of nlp approaches to modeling depression in social media: a post-covid-19 outlook. IEEE Journal of Biomedical and Health Informatics, pp. 1–13. External Links: Document Cited by: §1.
  • A. Bucur, M. Zampieri, T. Ranasinghe, and F. Crestani (2026) A survey on multilingual mental disorders detection from social media data. In Proceedings of the 19th Conference of the European Chapter of the Association for Computational Linguistics (Volume 1: Long Papers), Cited by: §1.
  • A. Ciobotaru, M. V. Constantinescu, L. P. Dinu, and S. Dumitrescu (2022) RED v2: enhancing RED dataset for multi-label emotion detection. In Proceedings of the Thirteenth Language Resources and Evaluation Conference, N. Calzolari, F. Béchet, P. Blache, K. Choukri, C. Cieri, T. Declerck, S. Goggi, H. Isahara, B. Maegaard, J. Mariani, H. Mazo, J. Odijk, and S. Piperidis (Eds.), Marseille, France, pp. 1392–1399. External Links: Link Cited by: §1, §4.
  • A. Ciobotaru and L. P. Dinu (2021) RED: a novel dataset for romanian emotion detection from tweets. In Proceedings of the international conference on recent advances in natural language processing (RANLP 2021), pp. 291–300. Cited by: §1.
  • Y. Dai, M. Jayaratne, and B. Jayatilleke (2022) Explainable personality prediction using answers to open-ended interview questions. 13. External Links: Document, Link, ISSN 1664-1078 Cited by: §2.
  • D. P. Dudău and F. A. Sava (2022) The development and validation of the romanian version of linguistic inquiry and word count 2015 (ro-liwc2015). Current Psychology 41 (6), pp. 3597–3614. Cited by: §1, §4.
  • S. Dumitrescu, A. Avram, and S. Pyysalo (2020) The birth of Romanian BERT. In Findings of the Association for Computational Linguistics: EMNLP 2020, T. Cohn, Y. He, and Y. Liu (Eds.), Online, pp. 4324–4328. External Links: Link, Document Cited by: §4.
  • S. Fekete (2002) The internet-a new source of data on suicide, depression and anxiety: a preliminary study. Archives of Suicide Research 6 (4), pp. 351–361. Cited by: §4.
  • L. Grillo (2016) A possible role of anhedonia as common substrate for depression and anxiety. 2016, pp. 1598130. External Links: Document Cited by: §5.2.
  • J. Hoying, B. M. Melnyk, E. Hutson, and A. Tan (2020) Prevalence and correlates of depression, anxiety, stress, healthy beliefs, and lifestyle behaviors in first-year graduate health sciences students. 17 (1), pp. 49–59. Note: Epub 2020 Jan 8 External Links: Document, Link Cited by: §5.1.
  • J. K. Hur, J. Heffner, G. W. Feng, J. Joormann, and R. B. Rutledge (2024) Language sentiment predicts changes in depressive symptoms. 121 (39), pp. e2321321121. Cited by: §4.
  • N. C. Jacobson and M. G. Newman (2014) Avoidance mediates the relationship between anxiety and depression over a decade later. 28 (5), pp. 437–445. Note: Epub 2014 Apr 16 External Links: Document, Link Cited by: §5.1.
  • N. H. Kalin (2020) The critical relationship between anxiety and depression. 177 (5), pp. 365–367. Note: PMID: 32354270 External Links: Document, Link, https://doi.org/10.1176/appi.ajp.2020.20030305 Cited by: §5.1.
  • G. Ke, Q. Meng, T. Finley, T. Wang, W. Chen, W. Ma, Q. Ye, and T. Liu (2017) Lightgbm: a highly efficient gradient boosting decision tree. Advances in neural information processing systems 30. Cited by: §4.
  • O. N. E. Kjell, P. Johnsson, and S. Sikström (2021) Freely generated word responses analyzed with artificial intelligence predict self-reported symptoms of depression, anxiety, and worry. 12, pp. 602581. External Links: Document, Link Cited by: §2.
  • K. Kroenke, R. L. Spitzer, and J. B. W. Williams (2001) The phq-9: validity of a brief depression severity measure. Journal of General Internal Medicine 16 (9), pp. 606–613. External Links: Document Cited by: §3, §4.
  • T. Liu, L. H. Ungar, B. Curtis, G. Sherman, K. Yadeta, L. Tay, J. C. Eichstaedt, and S. C. Guntuku (2022) Head versus heart: social media reveals differential language of loneliness from depression. 1 (1), pp. 16. Cited by: §1, §4.
  • D. M. Low, L. Rumker, T. Talkar, J. Torous, G. Cecchi, and S. S. Ghosh (2020) Natural language processing reveals vulnerable mental health support groups and heightened health anxiety on reddit during covid-19: observational study. 22 (10), pp. e22635. External Links: Document Cited by: §6.
  • S. M. Lundberg and S. Lee (2017) A unified approach to interpreting model predictions. Advances in neural information processing systems 30. Cited by: §4, §5.2.
  • S. H. Muhammad, N. Ousidhoum, I. Abdulmumin, J. P. Wahle, T. Ruas, et al. (2025) BRIGHTER: BRIdging the gap in human-annotated textual emotion recognition datasets for 28 languages. In Proceedings of the 63rd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), W. Che, J. Nabende, E. Shutova, and M. T. Pilehvar (Eds.), Vienna, Austria, pp. 8895–8916. External Links: Link, Document, ISBN 979-8-89176-251-0 Cited by: §1.
  • S. Nolen-Hoeksema, B. E. Wisco, and S. Lyubomirsky (2008) Rethinking rumination. Perspectives on psychological science 3 (5), pp. 400–424. Cited by: §5.2.
  • B. W. Penninx, Y. Milaneschi, F. Lamers, and N. Vogelzangs (2013) Understanding the somatic consequences of depression: biological mechanisms and the role of depression symptom profile. BMC medicine 11 (1), pp. 129. Cited by: §5.2.
  • S. S. C. Puspo, A. Bucur, S. Chancellor, O. Uzuner, and M. Zampieri (2026) Mental health disorder detection beyond social media: a systematic review of available datasets. In Proceedings of LREC, Cited by: §1.
  • A. B. S. Rahman, H. Ta, L. Najjar, A. Azadmanesh, and A. S. Gönül (2024) DepressionEmo: a novel dataset for multilabel classification of depression emotions. Note: Preprint External Links: Link Cited by: §2.
  • S. Rai, E. C. Stade, S. Giorgi, A. Francisco, L. H. Ungar, B. Curtis, and S. C. Guntuku (2024) Key language markers of depression on social media depend on race. Proceedings of the National Academy of Sciences 121 (14), pp. e2319837121. Cited by: §1, §1.
  • S. Rude, E. Gortner, and J. Pennebaker (2004) Language use of depressed and depression-vulnerable college students. Cognition & Emotion 18 (8), pp. 1121–1133. Cited by: §1, §4, §4.
  • A. H. Salopek and M. S. Eastin (2024) Toxic positivity intentions: an image management approach to upward social comparison and false self-presentation. Journal of Computer-Mediated Communication 29 (3), pp. zmae003. Cited by: §1.
  • S. Sikström, B. Kelmendi, and N. Persson (2023) Assessment of depression and anxiety in young and old with a question-based computational language approach. 2, pp. 11. External Links: Document, Link Cited by: §2.
  • R. L. Spitzer, K. Kroenke, J. B. W. Williams, and B. Löwe (2006) A brief measure for assessing generalized anxiety disorder: the gad-7. Archives of Internal Medicine 166 (10), pp. 1092–1097. External Links: Document Cited by: §3, §4.
  • R. N. Trifu, B. Nemeș, D. C. Herta, C. Bodea-Hategan, D. A. Talaș, and H. Coman (2024) Linguistic markers for major depressive disorder: a cross-sectional study using an automated procedure. Frontiers in Psychology 15, pp. 1355734. Cited by: §1, §1, §4, §5.2, §5.2, §6.
  • I. Van der Vegt and B. Kleinberg (2023) A multi-modal panel dataset to understand the psychological impact of the pandemic. Scientific Data 10 (1), pp. 537. External Links: Document, Link, ISSN 2052-4463 Cited by: §2.
  • A. Varastehnezhad, R. Tavasoli, S. Elyasi, M. LotfiNia, and H. Farbeh (2025) AI in mental health: emotional and sentiment analysis of large language models’ responses to depression, anxiety, and stress queries. abs/2508.11285. External Links: Link Cited by: §6.
  • S. Yoon, J. Jang, G. Son, S. Park, J. Hwang, J. Y. Choeh, and K. Choi (2024) Predicting neuroticism with open-ended response using natural language processing. Frontiers in Psychiatry 15, pp. 1437569. External Links: Document, Link Cited by: §2.
  • T. Zhang, K. Yang, S. Ji, and S. Ananiadou (2023) Emotion fusion for mental illness detection from social media: a survey. 92, pp. 231–246. External Links: ISSN 1566-2535, Document, Link Cited by: §2.